The Find two consecutive odd integers whose sum is 0
Answer by Earlsdon(6103)

Let the first odd integer be x and the next consecutive odd integer is x+2.
x + x+2 = 0
2x + 2 = 0
2x = -2
x = -1
x+2 = 1
The integers are: -1 and 1.